# Decision Synthesis in Monetary Policy

### v1

@@ Introduction
The BPDS forecast distributions are evaluated using importance sampling. At any given $\mathbf{x},$ the BPDS predictive distribution in eqn. (1) is simulated by sampling from the $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ in proportions defined by the BPDS probabilities $\pi_{j}(\mathbf{x})$; then the resulting importance sampling weights are simply proportional to the realized values of $\alpha_{j}(\mathbf{y}|\mathbf{x}).$ This provides for efficient computation as well as access to traditional methods and metrics– such as importance sampling effective sample sizes (ESS, e.g. Gruber and West, 2016, 2017, in related contexts)– to monitor and evaluate the quality of the resulting Monte Carlo approximations to resulting predictive expectations. Note that this can deliver such metrics to assess “concordance” between the initial densities $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ and their corresponding BPDS-tilted versions $f_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ in eqn. (2), as well as that of the initial mixture $p(\mathbf{y}|\mathbf{x})$ and the resulting $f(\mathbf{y}|\mathbf{x}).$ More aggressive BPDS target scores will generally lead to lower concordance, and choices can be partly guided by such empirical evaluations.
@@
The analysis and all empirical results proceed sequentially on an expanding window of data beginning in 1992Q2. Our summaries begin with a comparison of the decisions recommended by BPDS to those suggested by BMA. This is followed by a discussion of the individual models and how they are combined by BPDS and BMA. Additional discussion highlights some operational BPDS details to provide further insights into the resulting decision outcomes.
@@
Figure 1 shows the actual policy rate each quarter along with the $1{-}8$ quarters ahead policy recommendations that would have been made by BPDS and BMA. In using the shadow rate, the zero lower bound is not in effect and negative values for the policy rate are possible. Recommendations for negative values for the policy rate are not to be taken literally as advising cuts to a negative Federal Funds rate, but rather a suggestion to undertake other forms of monetary easing that would be expected to proxy such cuts.
Since 2014 the optimal policy paths recommended by the two approaches are generally similar, though there are notable differences prior to that time. Some specific periods of interest are now highlighted.
@@
During this period, BMA and BPDS provide similar recommendations that are often quite different from the actual policy rate. For almost all of these times, the policy recommendations are to cut interest rates whereas (apart from 2019-2021) the actual policy rate increased. Some differences do arise between BMA and BPDS, for example, during the post-COVID inflation BPDS recommends a higher rate path. BPDS is closer to the decision actually made by the Federal Reserve, although according to BPDS rate cuts should have begun coming down by now.
@@
It is during this period that the differences between BPDS and BMA are most acute. The actual policy rate fell slowly during this period. BPDS recommends rate cuts as well, initially at a more rapid rate than what actually occurred, but as of 2010, its recommendations are similar to the ones the policy makers actually made. In contrast, BMA recommends huge cuts to the policy rate right at the start of the financial crisis but subsequently consistently argues for rate increases.
@@
In the period from 2003 through the beginning of the financial crisis, the actual policy rate was gradually increasing. In this period, BMA consistently recommended rapid rate increases. In contrast, BPDS recommendations were generally similar to what actually transpired, apart from at the beginning of this period where the advice was to raise the policy rate more slowly than occurred.
@@
During this period, the pattern is more mixed. Optimal policy recommendations generated under each of BMA and BPDS often differed from actual decisions, with no consistent pattern; at times the recommended rates were higher than the actual policy rate, and at other times they were lower.
@@
A general pattern, one that occurs throughout the sample period, is that BMA and BPDS typically recommend larger changes in policy rates than were actually implemented by policy makers. Part of this is presumably due to differences between the policy makers utility function and those used in our analyses. Other possible explanations are that policy makers can affect expectations through their communications which is a channel not captured in the model or their models use a much steeper Phillips Curve. Also, we focus only on inflation up to 2 years ahead, without considering the possibility of an over- or under-shoot of inflation after 8 quarters while policy makers would generally desire longer-term targets sustainably.
@@
Figures 24 shed light on these patterns. These images represent the time trajectories of predictive densities of inflation at each relevant horizon, using BPDS (Fig. 2) and BMA (Fig. 3), as well as their differences (Fig. 4). These indicate that the BPDS mixtures are less dispersed than the BMA mixture for much of the sample period, i.e., BMA predictive distributions are relatively more heavy-tailed, especially at longer horizons. This is partly due to the BPDS score function emphasizing that the policy maker wants to avoid extreme inflation outcomes, and also accounts for why BMA often tells policy makers to make larger changes to the policy rate than BPDS as discussed in Section 4.2. The differences between BPDS and BMA become larger at longer forecast horizons. Medium and long-term macroeconomic forecasting is difficult which leads to standard methods such as BMA producing fairly dispersed predictive densities at longer horizons. BPDS, on the other hand, is reducing this effect, which dampens the BPDS optimal decisions and reduces predictive uncertainty relative to BMA. Then, differences between BPDS and BMA forecast densities are reduced after the financial crisis, which helps account for why their policy recommendations are similar in the last decade of the sample.
Figure 2: BPDS forecast densities of inflation. The frames represent 1–8 quarter ahead forecasts, reading along the rows from top-left to bottom-right. The colors represent probabilities with the shading from blue, being lower probability, to red being higher probability
Figure 3: BMA forecast densities of inflation. The frames represent 1–8 quarter ahead forecasts, reading along the rows from top-left to bottom-right. The colors represent probabilities with the shading from blue, being lower probability, to red being higher probability
Figure 4: Difference between BPDS and BMA forecast densities of inflation. Red-shaded regions have higher probability under BPDS than under BMA, with blue shading indicating the reverse, and white shading equal probability
@@
Figure 5 shows the trajectories of model probabilities under BPDS and BMA. These are the discounted AVS prior model probabilities $\pi_{tj}$ over time, the implied initial decision-dependent probabilities $\pi_{tj}(\mathbf{x}_{t})$ evaluated at the BPDS-optimal decision $\mathbf{x}_{t}$ at each time, the resulting BPDS probabilities $\tilde{\pi}_{tj}(\mathbf{x}_{t})$ of eqn. (3) at each time, and the standard BMA probabilities over time.
@@
Under traditional BMA, the two model probabilities are appreciable until the financial crisis. After the start of the crisis, the less parsimonious ${\mathcal{M}}_{2}$– that includes additional financial variables– receives virtually all the weight. In contrast, BPDS weights vary more over time, allocating most of the weight to the parsimonious ${\mathcal{M}}_{1}$ for much of the period (i.e. 1997 through 2017), though ${\mathcal{M}}_{2}$ plays more of a role at both the beginning and end of the sample period. That BPDS generally favors the more parsimonious ${\mathcal{M}}_{1}$, with less dispersed forecast distributions, partially accounts for why BPDS often dampens extreme recommendations made when using BMA.
BPDS probabilities on the over-dispersed ${\mathcal{M}}_{0}$ are generally small, though with notable increases at two critical periods: the start of the financial crisis and the start of the COVID-19 pandemic. In such extreme times, when neither ${\mathcal{M}}_{1}$ or ${\mathcal{M}}_{2}$ forecasts well, the increased probability on the fall-back ${\mathcal{M}}_{0}$– though small– provides an indicator of this.
The BPDS prior model probabilities $\pi_{tj}$ based on discounted AVS differ noticeably from BMA probabilities (except at the very start of the time period). A big impact then arises from the conditioning on information generated in the decision space to map these $\pi_{tj}$ to the decision-dependent weights $\pi_{tj}(\mathbf{x}_{t})$ at the BPDS optimal decisions $\mathbf{x}_{t}$ at each time. Recall that this mapping theoretically properly takes into account the likelihood of future, as yet unobserved, interest rate outcomes; this relevant information is not accounted for in the prior weights $\pi_{tj}$ and is, of course, absent under BMA. The subsequent map from initial probabilities $\pi_{tj}(\mathbf{x}_{t})$ to the BPDS weights $\tilde{\pi}_{tj}(\mathbf{x}_{t})$ is wholly based on the impact of the entropic tilting towards “more favorable” decisions, in expectation. We see that the impact is rather small over time, and this is to be expected: the BPDS analysis uses “small” perturbations of the initial mixture based on target expected scores that are only modest increases over those under the initial mixture. We expect to see slight tilting towards models that are expected to do well, but not large changes relative to the initial probabilities.
@@
Time trajectories of the evaluated tilting vectors ${\bm{\tau}}_{t}(\mathbf{x}_{t})$, evaluated at the optimal decisions $\mathbf{x}_{t}$, are shown in Figure 7. The values generally tend to increase with horizon $h$, thus attaching more weight to longer forecasting horizons. This is partly to be expected due to the higher uncertainties at longer forecast horizons.
@@
Figure 6: Trajectories of the 8 elements of the evaluated BPDS tilting vector ${\bm{\tau}}_{t}(\mathbf{x}_{t})$ at the optimized $\mathbf{x}_{t}$ at each quarter
@@
Figure 7: Trajectories of the effective sample size (ESS) metrics for individual models and for the BPDS initial mixture
@@
Figure 7 plots trajectories of several effective sample size (ESS) measures arising from the importance sampling to simulate BPDS predictive distributions as discussed in Section 3.3. This provides a read-out of the extent of tilting the initial mixture $p(\mathbf{y}|\mathbf{x})$ to the BPDS mixture $f(\mathbf{y}|\mathbf{x})$ as well as that for tilting each of the individual model pdfs from $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ to $f_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ (again, time-indexed and updated throughout the time series). Until the COVID recession, the ESS of the initial mixture is stable between 90-95% suggesting only a small amount of tilting, as desired. The COVID recession is a period of rapid change as expected, as we see large changes in the initial weights and larger values of ${\bm{\tau}}$ required to achieve the desired target. The low value of ESS indicates that at that point the target expected scores are unrealistic given the then current state of the economy. However, the resulting decisions during this time period appear to be rather sensible, meaning we do not need to be too concerned about the low ESS which can, in any case, be redressed by simply increasing the overall Monte Carlo sample size accordingly. The ESS values of individual models are generally lower than that of the overall mixture and somewhat more volatile. One nice point is that, even when one of the models seems to suffer a low ESS, the BPDS mixture ESS is generally maintained at higher values. This indicates that BPDS is able to strike a balance in weighting expected versus historical performance of models on both predictive and decision outcomes.
@@
Finally, Figure 8 compares the realized trajectories of expected utilities under BPDS and BMA. Each uses the same utility function to define the final optimal policy path decision, so these are directly comparable, and the comparison is relevant in terms of the setting of forward, sequential decisions where a change to much lower values at any time point should signal concern to the decision-maker. BPDS is designed to target an expected utility higher than that of the initial mixture, but whether it achieves a higher expected utility than BMA– which has different initial probabilities and lacks outcome-dependent weighting– is a question for empirical study. In this example, as illustrated in the figure, BPDS utility does exceed that of BMA in virtually every period. After the financial crisis, the two are similar, consistent with the earlier finding that they typically produced similar decisions during this time period. However, before the financial crisis, there were several periods during which the BPDS expected utilities were substantially higher than those of BMA. These correspond to times where we see more differences between optimal policy path recommendations, and within which there are some periods of greater concordance between BPDS and actual policy decisions as well as more constrained (i.e., less extreme) recommended decisions under BPDS relative to BMA.
Figure 8: Trajectories of expected utilities comparing BPDS with BMA
@@
Earlier applications of BPDS have focused mainly on financial portfolio forecasting and decisions (Tallman and West, 2023, section 6; Tallman and West, 2024), a setting in which forecasting models do not (generally) depend on the decisions of interest, while utility functions may and often do depend on the models and their predictions. In contrast, the setting of monetary policy analysis is one in which the dependence of models and their forecasts on the decision variables (policy instruments) is simply fundamental. It is also a setting in which the decision variables are treated simultaneously as outcomes. The future paths of central bank interest rates, for example, are modelled as time series outcomes along with other economic and financial indicators in VAR models. This then leads to conditioning on decision variables to define predictions of other indicators, with consequent implications for relative model weights in the model uncertainty setting. This latter point is critical as it then leads to relatively up- or down-weighting a model based on how well-supported a particular candidate decision is under its predictions; to our knowledge, this is the first time this central question has been formally, statistically addressed. These central features of predictive decision-making in monetary policy contexts are addressed with extensions and customization of the existing theory of BPDS.
@@
The BPDS perspective– of integrating historical and expected decision outcomes with focused aspects of statistical predictive performance into relative model weightings– is new to the policy arena. We argue for this perspective since policy makers are primarily interested in utilizing sets of models for the eventual policy decisions. Pure forecasting exercises– and evaluation and combinations of models for prediction per se– is, of course, of parallel interest and importance. We emphasize that BPDS also involves addressing predictive performance on specific, defined outcomes of interest. But most importantly, by putting the spotlight on decision-making, we gain additional insights into policy making that are not possible in exercises that focus solely on predictive performance.
@@
In a recursive real-time decision-making exercise, we find substantial differences at various periods of time between the policy recommendations of BPDS and the traditional Bayesian model averaging approach, though good concordance at other times. When recommended policy decisions differ between the approaches, in most cases the BPDS policy paths are more intuitively sensible and less extreme than under BMA, and more consistent with the actual decisions made by the policy makers at the time. The case study presented investigates and interprets aspects of BPDS– in terms of differential model weights based on historical information alone and then updated based on identified optimal decisions, with consequent insights into how the differences relative to standard BMA arise and are exploited. This case study is a first step towards broader development and evaluation of BPDS in a setting with larger numbers of econometric models. The parallel next steps will naturally include BPDS for scenario forecasting– analyses in collaboration with policy-makers that explore and aim to understand the sensitivity of model-based recommendations relative to chosen potential economic scenarios.

### v2

@@ Introduction
BPDS forecast distributions are evaluated using importance sampling. At any given $\mathbf{x},$ the BPDS predictive distribution in eqn. (1) is simulated by sampling from the $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ in proportions defined by the BPDS probabilities $\pi_{j}(\mathbf{x})$. Importance sampling weights are then proportional to the realized values of $\alpha_{j}(\mathbf{y}|\mathbf{x}).$ This provides for efficient computation as well as access to traditional methods and metrics– such as the importance of sampling effective sample sizes (ESS– e.g., Gruber and West, 2016, 2017, in related contexts)– to monitor and evaluate the quality of the resulting Monte Carlo approximations compared to resulting predictive expectations. Note that this evaluation can deliver such metrics to assess “concordance” between the initial densities $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ and their corresponding BPDS-tilted versions $f_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ in eqn. (2), as well as that of the initial mixture $p(\mathbf{y}|\mathbf{x})$ and the resulting $f(\mathbf{y}|\mathbf{x}).$ More aggressive BPDS target scores will generally lead to lower concordance, and choices can be partly guided by such empirical evaluations.
@@
Analyses proceed sequentially on an expanding window of data beginning in 1992Q2. We compare decisions recommended by BPDS to those under BMA, discuss the individual models and how they are combined by BPDS and BMA, and highlight some operational BPDS details underlying insights into the resulting decision outcomes.
@@
Figure 1 shows the actual policy rate each quarter along with the $1{-}8$-quarters-ahead policy recommendations that would have been made by BPDS and BMA. In using the shadow rate, the zero lower bound is not in effect and negative values for the policy rate are possible. Recommendations for negative values for the policy rate are not to be taken literally as advising cuts to a negative Federal Funds rate, but rather as a suggestion to undertake other forms of monetary easing that would be expected to proxy such cuts.
Since 2014, the optimal policy paths recommended by the two approaches are generally similar, though there are notable differences prior to that time. Some specific periods of interest are now highlighted.
@@
During this period, BMA and BPDS provide similar recommendations that are often quite different from the actual policy rate. For almost all of these times, the policy recommendations are to cut interest rates, whereas (apart from 2019–2021) the actual policy rate increased. Some differences do arise between BMA and BPDS. For example, during the post-COVID inflation period, BPDS recommends a higher rate path. BPDS is closer to the decision actually made by the Federal Reserve, although according to BPDS, interest rates should decrease throughout 2023 and 2024.
@@
It is during this period that the differences between BPDS and BMA are most acute. The actual policy rate fell slowly during this period. BPDS recommends rate cuts as well, initially at a more rapid rate than what actually occurred, but as of 2010, its recommendations are similar to the ones the policymakers actually made. In contrast, BMA recommends huge cuts to the policy rate right at the start of the financial crisis, but subsequently consistently argues for rate increases.
@@
From 2003 through to the beginning of the financial crisis, the actual policy rate was gradually increasing. In this period, BMA consistently recommends rapid rate increases. In contrast, BPDS recommendations are generally similar to what actually transpired, apart from at the beginning of this period where the advice is to raise the policy rate more slowly than what actually occurred.
@@
During this period, the pattern is more mixed. Optimal policy recommendations generated under each of BMA and BPDS often differ from actual decisions, with no consistent pattern; at times the recommended rates are higher than the actual policy rate, and at other times they are lower.
@@
A general pattern, one that occurs throughout the sample period, is that BMA and BPDS typically recommend larger changes in policy rates than were actually implemented by policymakers. Part of this is presumably due to differences between the policymakers’ utility function and those used in our analyses. Other possible explanations are that policymakers can affect expectations through their communications, which is a channel not captured in the model, or their models use a much steeper Phillips Curve. Also, we focus only on inflation up to two years ahead, without considering the possibility of an over- or under-shoot of inflation after eight quarters. In contrast, policymakers would generally aim for inflation to be sustainably at target over the longer-term.
@@
Figures 24 shed light on these patterns. These images represent the time trajectories of predictive densities of inflation at each relevant horizon, using BPDS (Fig. 2) and BMA (Fig. 3), as well as their differences (Fig. 4). These indicate that the BPDS mixtures are less dispersed than the BMA mixture for much of the sample period; that is, BMA predictive distributions are relatively more heavy-tailed, especially at longer horizons. This is partly due to the BPDS score function emphasizing that the policymaker wants to avoid extreme inflation outcomes, and also accounts for why BMA often tells policymakers to make larger changes to the policy rate than BPDS, as discussed in Section 4.2. The differences between BPDS and BMA become larger at longer forecast horizons. Medium- and long-term macroeconomic forecasting is difficult, which leads to standard methods such as BMA producing fairly dispersed predictive densities at longer horizons. BPDS, on the other hand, is reducing this effect, which dampens the BPDS optimal decisions and reduces predictive uncertainty relative to BMA. Then, differences between BPDS and BMA forecast densities are reduced after the financial crisis, which helps account for why their policy recommendations are similar in the last decade of the sample.
Figure 2: BPDS forecast densities of inflation. The frames represent 1–8 quarter ahead forecasts, reading along the rows from top-left to bottom-right. The colours represent probabilities, with the blue shading showing lower probability, and red showing higher probability.
Figure 3: BMA forecast densities of inflation. The frames represent 1–8 quarter ahead forecasts, reading along the rows from top-left to bottom-right. The colours represent probabilities, with the blue shading showing lower probability, and red showing higher probability.
Figure 4: Difference between BPDS and BMA forecast densities of inflation. Red-shaded regions have higher probability under BPDS than under BMA, with blue shading indicating the reverse, and white shading showing equal probability.
@@
Figure 5 shows trajectories of model probabilities under BPDS and BMA. At each $t$, these are the discounted AVS prior model probabilities $\pi_{tj}$, implied initial decision-dependent probabilities $\pi_{tj}(\mathbf{x}_{t})$ evaluated at the BPDS-optimal decision $\mathbf{x}_{t}$, resulting BPDS probabilities $\tilde{\pi}_{tj}(\mathbf{x}_{t})$ of eqn. (3), and standard BMA probabilities.
@@
Under traditional BMA, the two model probabilities are appreciable until the financial crisis. After the start of the crisis, the less parsimonious ${\mathcal{M}}_{2}$, which includes additional financial variables, receives virtually all the weight. In contrast, BPDS weights vary more over time, allocating most of the weight to the parsimonious ${\mathcal{M}}_{1}$ for much of the period (i.e., 1997 through 2017), though ${\mathcal{M}}_{2}$ plays more of a role at both the beginning and end of the sample period. That BPDS generally favours the more parsimonious ${\mathcal{M}}_{1}$, with less dispersed forecast distributions, partially accounts for why BPDS often dampens extreme recommendations made when using BMA.
BPDS probabilities on the over-dispersed ${\mathcal{M}}_{0}$ are generally small, though with notable increases at the start of the COVID-19 pandemic. In such extreme times, when neither ${\mathcal{M}}_{1}$ nor ${\mathcal{M}}_{2}$ forecasts well, the increased probability on the fall-back ${\mathcal{M}}_{0}$– though small– provides an indicator of this.
The BPDS prior model probabilities $\pi_{tj}$ based on discounted AVS differ noticeably from BMA probabilities (except at the very start of the time period). A big impact then arises from the conditioning on information generated in the decision space to map these $\pi_{tj}$ to the decision-dependent weights $\pi_{tj}(\mathbf{x}_{t})$ at the BPDS optimal decisions $\mathbf{x}_{t}$ at each time. Recall that this mapping theoretically properly takes into account the likelihood of future, as yet unobserved, interest rate outcomes; this relevant information is not accounted for in the prior weights $\pi_{tj}$ and is, of course, absent under BMA. The subsequent map from initial probabilities $\pi_{tj}(\mathbf{x}_{t})$ to the BPDS weights $\tilde{\pi}_{tj}(\mathbf{x}_{t})$ is wholly based on the impact of the entropic tilting towards “more favourable” decisions in expectation. We see that the impact is rather small over time, and this is to be expected: the BPDS analysis uses “small” perturbations of the initial mixture based on target expected scores that are only modest increases over those under the initial mixture. We expect to see slight tilting towards models that are expected to do well, but not large changes relative to the initial probabilities.
@@
Time trajectories of the tilting vectors ${\bm{\tau}}_{t}(\mathbf{x}_{t})$ evaluated at the optimal decisions $\mathbf{x}_{t}$ are shown in Figure 7. The values generally tend to increase with horizon $h$, thus attaching more weight to longer forecasting horizons. This is partly to be expected due to the higher uncertainties at longer forecast horizons.
@@
Figure 6: Trajectories of the sixteen elements of the evaluated BPDS tilting vector ${\bm{\tau}}_{t}(\mathbf{x}_{t})$ at the optimized $\mathbf{x}_{t}$ at each quarter. (a) ${\bm{\tau}}_{t}(\mathbf{x}_{t})$ associated with inflation (b) ${\bm{\tau}}_{t}(\mathbf{x}_{t})$ associated with GDP.
@@
Figure 7: Trajectories of the effective sample size (ESS) metrics for individual models and for the BPDS initial mixture.
@@
Figure 7 plots trajectories of several effective sample size (ESS) measures arising from the importance sampling to simulate BPDS predictive distributions, as discussed in Section 3.3. This provides a read-out of the extent of tilting the initial mixture $p(\mathbf{y}|\mathbf{x})$ to the BPDS mixture $f(\mathbf{y}|\mathbf{x})$, as well as that for tilting each of the individual model pdfs from $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ to $f_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j})$ (again, time-indexed and updated throughout the time series). Until the COVID recession, the ESS of the initial mixture is stable between 90–95%, suggesting only a small amount of tilting, as desired. The COVID recession is a period of rapid change, as expected, as we see large changes in the initial weights and larger values of ${\bm{\tau}}$ required to achieve the desired target. The low value of ESS indicates that at that point, the target expected scores are unrealistic given the then-current state of the economy. However, the resulting decisions during this time period appear to be rather sensible. This means we do not need to be too concerned about the low ESS, which can, in any case, be redressed by simply increasing the overall Monte Carlo sample size accordingly. The ESS values of individual models are generally lower than that of the overall mixture and somewhat more volatile. One nice point is that, even when one of the models seems to suffer a low ESS, the BPDS mixture ESS is generally maintained at higher values. This indicates that BPDS is able to strike a balance in weighting expected versus historical performance of models on both predictive and decision outcomes.
@@
Finally, Figure 8 compares the realized trajectories of expected utilities under BPDS and BMA. Each uses the same utility function to define the final optimal policy path decision, so these are directly comparable, and the comparison is relevant in terms of the setting of forward, sequential decisions where a change to much lower values at any time point should signal concern to the decision-maker. BPDS is designed to target an expected utility higher than that of the initial mixture, but whether it achieves a higher expected utility than BMA– which has different initial probabilities and lacks outcome-dependent weighting– is a question for empirical study. In this example, as illustrated in the figure, BPDS utility does exceed that of BMA in virtually every period. After the financial crisis, the two are similar, consistent with the earlier finding that they typically produce similar decisions during this time period. However, before the financial crisis, there are several periods during which the BPDS expected utilities are substantially higher than those of BMA. These correspond to times where we see more differences between optimal policy path recommendations. Within these times there are some periods of greater concordance between BPDS and actual policy decisions, as well as more constrained (i.e., less extreme) recommended decisions under BPDS relative to BMA.
Figure 8: Trajectories of expected utilities comparing BPDS with BMA.
@@
Earlier applications of BPDS focused on financial forecasting and portfolio decisions (Tallman and West, 2023, 2025). In this setting, forecasting models do not (generally) depend on the decisions of interest, while utility functions may and often do depend on the models and their predictions. In contrast, the setting of monetary policy analysis is one in which the dependence of models and their forecasts on the decision variables (policy instruments) is simply fundamental. It is also a setting in which the decision variables are treated simultaneously as outcomes. The future paths of central bank interest rates, for example, are modelled as time series outcomes along with other economic and financial indicators in VAR models. This then leads to conditioning on decision variables to define predictions of other indicators, with consequent implications for relative model weights in the model uncertainty setting. This latter point is critical as it then leads to relatively up- or down-weighting a model based on how well-supported a particular candidate decision is under its predictions; to our knowledge, this is the first time this central question has been formally, statistically addressed. These central features of predictive decision-making in monetary policy contexts are addressed with extensions and customization of the existing theory of BPDS.
@@
The BPDS perspective– of integrating historical and expected decision outcomes with focused aspects of statistical predictive performance into relative model weightings– is new to the policy arena. We argue for this perspective since policymakers are primarily interested in using sets of models for the eventual policy decisions. Pure forecasting exercises– and evaluation and combinations of models for prediction per se– are, of course, of parallel interest and importance. We emphasize that BPDS also involves addressing predictive performance on specific, defined outcomes of interest. But most importantly, by putting the spotlight on decision-making, we gain additional insights into policy-making that are not possible in exercises that focus solely on predictive performance.
@@
In a recursive, real-time decision-making exercise, we find substantial differences at various periods of time between the policy recommendations of BPDS and the traditional Bayesian model averaging approach, though good concordance at other times. When recommended policy decisions differ between the approaches, in most cases the BPDS policy paths are more intuitively sensible and less extreme than under BMA, and more consistent with the actual decisions made by the policymakers at the time. The case study presented investigates and interprets aspects of BPDS in terms of differential model weights based on historical information alone, and then updated based on identified optimal decisions, with consequent insights into how the differences relative to standard BMA arise and are exploited. This case study is a first step towards broader development and evaluation of BPDS in a setting with larger numbers of econometric models. Parallel next steps can naturally be expected include broader evaluation in terms of multiple models, and developments for scenario forecasting. Such extensions, in collaboration with policymakers, will advance understanding the sensitivity of model-based recommendations relative to chosen potential economic scenarios.